# A new model of spontaneous scalarization of black holes induced by curvature and matter

### v1

@@ III.3 Thermodynamic properties
The black hole entropy $S_{H}$ in this model presents corrections to the Bekenstein-Hawking law, due to the non-minimal coupling with the Gauss-Bonnet term. We adopt the Wald–Iyer approach [151, 152], which associates a Noether charge to the black hole entropy in theories with diffeomorphism-invariant Lagrangian: then the entropy $S_{H}$ is defined as
@@
For instance scalarized black holes in EMsGB theories satisfy Smarr relation [152, 153, 86]. For the static solutions considered here ($J=0,\Omega_{H}=0$), this relation is:
@@
where $M_{s}$ is a contribution of the scalar field [154]:
@@
The most useful quantities to plot for studying black hole thermodynamic properties are the horizon area, the Hawking temperature, and the entropy. The entropy can be plotted as $S_{H}/\eta$ or as $S_{H}/S_{GR}$. The temperature and area can be represented either normalizing them with respect to the coupling constant, for instance $A_{H}/\eta$ and $T_{H}/\eta$, or by using the reduced quantities :
@@
Figure 4: The figures present several plots of the reduced area $a_{H}=\frac{A_{H}}{16\pi M^{2}}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each of them shows many solutions with different values of $\alpha$.
In Fig. 4 is reported the reduced area with respect to the reduced charge for several values of the other parameters: we can notice that the first two plots, corresponding to $Q/\sqrt{\eta}=0.2$ and 0.5473 show that the area of all the scalarized solutions is smaller than the RN one, while in the third plot, with $Q/\sqrt{\eta}=0.8$, the behavior of the solutions with $\alpha=-20,-30,-40$ is opposite, showing scalarized black holes with higher area than the RN one, which is in agreement with the fact that matter-induced scalarization becomes dominant due to the high value of the coupling to Maxwell ($\alpha$). From the value $Q/\sqrt{\eta}=1.6$ we clearly see that all the scalarized solutions with different $\alpha$ show higher area with respect to the RN one: this is expected even for the smallest value of $\alpha$ because of the high value of $Q/\sqrt{\eta}$ which enhances the contribution of scalarization induced by Maxwell in comparison with the one induced by the GB invariant.
Figure 5: The figures present several plots of the reduced temperature $t_{H}=8\pi MT_{H}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each one shows many solutions with different values of $\alpha$.
In Fig. 5 the reduced temperature of the scalarized solutions is plotted against the reduced charge. One can see that all scalarized solutions bifurcate from the GR branch, and have higher temperature than the RN black hole. Another important property is that they never reach extremality ($t_{H}$ never vanishes). Finally, it turns out that, starting from $Q/\sqrt{\eta}=1.6$, the scalarized black hole can become overcharged ($q>1$): this fact was never observed for $GB^{+}$ ($\eta>0$) scalarization before, but only in the case of $GB^{-}$ ($\eta<0$) as pointed out in [86].
Figure 6: The figures show the entropy of scalarized black holes normalized by the entropy of GR black holes with the same parameters ($S_{H}/S_{GR}$), with respect to the normalized mass $M/\sqrt{\eta}$: each one shows many solutions with different values of $\alpha$.
In Fig. 6 is depicted the entropy of scalarized black holes with respect to their GR counterpart. We see clearly that the entropy of all scalarized black holes is larger than the entropy of RN black holes, so they are thermodynamically preferred. For large values of $Q/\sqrt{\eta}$, the entropy is almost the same as in GR, especially for larger masses $M/\sqrt{\eta}$, while for small values of $Q/\sqrt{\eta}$ the entropy of scalarized black holes deviates from its GR value.
@@ IV Conclusions
For different choices of $\alpha$ and $\eta$, the curves can follow different trajectories, but interestingly many branches tend towards similar values of scalar charge as they approach their endpoints.
By comparing the branches, we found indications of which Maxwell coupling may stabilize solutions with quadratic coupling as the Ricci coupling and self interactions did.
@@
Many future directions to extend this work are possible. One first step should derive the black hole solutions for $\eta<0$ ($GB^{-}$). The next one will be to study the stability of these solutions (radial and angular), the work of [102] make it simple to do the analysis. Future work can examine different coupling functions in our model. For instance, using an exponential form of coupling for one of the invariants or both of them instead of the quadratic forms.
@@
Our new model with both curvature and matter couplings broadens the domain of spontaneous black hole scalarization and reveals a much richer structure of solutions. We have shown that scalarized black holes in this theory can exist under more flexible conditions, they appear to be more stable, possess higher entropy, and even achieve charge-to-mass ratios beyond the traditional extremal limit. These findings deepen our understanding of scalarization phenomena by introducing a new type of scalarization induced by two different sources and set the stage for further investigations into the stability and astrophysical implications of scalarized black holes.
@@ Appendix A Field Equations of motion for the model
$\quad-2rf_{,\phi}(\phi)\left(r^{4}g(\phi)^{2}\left(2r^{2}\phi^{\prime 2}\right.\right.$
@@
$Z$ $=-2r^{2}g(\phi)N\left(8Q^{2}(N-1)f_{,\phi}(\phi)^{2}\left((3N+1)f_{,\phi}(\phi)\phi^{\prime}-r\right)+r^{2}g(\phi)\left(r^{5}+2f_{,\phi}(\phi)\left((1-7N)\phi^{\prime}r^{4}\right.\right.\right.$

### v2

@@ III.3 Thermodynamic properties
The black hole entropy $S_{H}$ in this model presents corrections to the Bekenstein-Hawking law, due to the non-minimal coupling with the Gauss-Bonnet term. We adopt the Wald–Iyer approach [155, 156], which associates a Noether charge to the black hole entropy in theories with diffeomorphism-invariant Lagrangian: Then, the entropy $S_{H}$ is defined as
@@
For instance scalarized black holes in EMsGB theories satisfy the Smarr relation [156, 157, 86]. For the static solutions considered here ($J=0,\Omega_{H}=0$), this relation is
@@
where $M_{s}$ is a contribution of the scalar field [158],
@@
The most useful quantities to plot for studying black hole thermodynamic properties are the horizon area, the Hawking temperature, and the entropy. The entropy can be plotted as $S_{H}/\eta$ or as $S_{H}/S_{GR}$. The temperature and area can be represented either by normalizing them with respect to the coupling constant, for instance $A_{H}/\eta$ and $T_{H}/\eta$, or by using the reduced quantities
@@
Figure 4: The figures present several plots of the reduced area $a_{H}=\frac{A_{H}}{16\pi M^{2}}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$, and each of them shows many solutions with different values of $\alpha$.
In Fig. 4 the reduced area is reported with respect to the reduced charge for several values of the other parameters. We can notice that the first two plots, corresponding to $Q/\sqrt{\eta}=0.2$ and 0.5473 show that the area of all the scalarized solutions is smaller than the RN one, while in the third plot, with $Q/\sqrt{\eta}=0.8$, the behavior of the solutions with $\alpha=-20,-30,-40$ is opposite, showing scalarized black holes with higher area than the RN one, which is in agreement with the fact that matter-induced scalarization becomes dominant due to the high value of the coupling to Maxwell ($\alpha$). From the value $Q/\sqrt{\eta}=1.6$ we clearly see that all the scalarized solutions with different values of $\alpha$ show higher area with respect to the RN one. This is expected even for the smallest value of $\alpha$ because of the high value of $Q/\sqrt{\eta}$ which enhances the contribution of scalarization induced by Maxwell in comparison with the one induced by the GB invariant.
Figure 5: The figures present several plots of the reduced temperature $t_{H}=8\pi MT_{H}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$, and each one shows many solutions with different values of $\alpha$.
In Fig. 5 the reduced temperature of the scalarized solutions is plotted against the reduced charge. One can see that all scalarized solutions bifurcate from the GR branch, and have higher temperatures than the RN black hole. Another important property is that they never reach extremality ($t_{H}$ never vanishes). Finally, it turns out that, starting from $Q/\sqrt{\eta}=1.6$, the scalarized black hole can become overcharged ($q>1$). This fact was never observed for $GB^{+}$ ($\eta>0$) scalarization before, but only in the case of $GB^{-}$ ($\eta<0$) as pointed out in [86].
Figure 6: The figures show the entropy of scalarized black holes normalized by the entropy of GR black holes with the same parameters ($S_{H}/S_{GR}$), with respect to the normalized mass $M/\sqrt{\eta}$, and each one shows many solutions with different values of $\alpha$.
In Fig. 6, the entropy of scalarized black holes is depicted with respect to their GR counterparts. We see clearly that the entropy of all scalarized black holes is larger than the entropy of RN black holes, so they are thermodynamically preferred. For large values of $Q/\sqrt{\eta}$, the entropy is almost the same as in GR, especially for larger masses $M/\sqrt{\eta}$, while for small values of $Q/\sqrt{\eta}$ the entropy of scalarized black holes deviates from its GR value.
@@ IV Conclusions
For different choices of $\alpha$ and $\eta$, the curves can follow different trajectories, but interestingly many branches tend toward similar values of scalar charge as they approach their end points.
By comparing the branches, we found indications of which Maxwell coupling may stabilize solutions with quadratic coupling as the Ricci coupling and self-interactions did.
@@
Many future directions to extend this work are possible. One first step should derive the black hole solutions for $\eta<0$ ($GB^{-}$). The next one will be to study the stability of these solutions (radial and angular), where the work of Ref [106] will make it simple to do the analysis. Future work can examine different coupling functions in our model. For instance, using an exponential form of coupling for one of the invariants or both of them instead of the quadratic forms.
@@
Our new model with both curvature and matter couplings broadens the domain of spontaneous black hole scalarization and reveals a much richer structure of solutions. We have shown that scalarized black holes in this theory can exist under more flexible conditions, where they appear to be more stable, possess higher entropy, and even achieve charge-to-mass ratios beyond the traditional extremal limit. These findings deepen our understanding of scalarization phenomena by introducing a new type of scalarization induced by two different sources and set the stage for further investigations into the stability and astrophysical implications of scalarized black holes.
@@ Appendix A Field equations of motion for the model
[…]
@@
[…]